OpenStudy (anonymous):

What is the solution set of |2x – 3| – 8 = –1? 7 or –7 5 or –2 5 or 2 No solution

OpenStudy (goformit100):

5 or 2

Parth (parthkohli):

Add 8 to both sides first. \( \color{Black}{\Rightarrow |2x - 3| = 7 }\) So, \(2x - 3 = \pm 7\)

Parth (parthkohli):

Two equations to solve here: \( \color{Black}{\Rightarrow 2x - 3 = 7 }\) \( \color{Black}{\Rightarrow 2 x - 3 = -7 }\)

OpenStudy (shubhamsrg):

why dont you go for option checking instead of actually solving!! (saves time) ;)

OpenStudy (lgbasallote):

^that depends lol

OpenStudy (sriram):

means 2x-3=7 or -7 hence solving it we get 2x=10 or -4 giving x=5 or -2

OpenStudy (lgbasallote):

some people can solve this faster than checking blindly ;)

Parth (parthkohli):

I think that understanding is the key factor and solving it is the best way to understand the problem. I prefer solving rather than checking.

OpenStudy (shubhamsrg):

hmmn.. :|